# Words Matter: Leveraging Individual Text Embeddings for Code Generation in CLIP Test-Time Adaptation

### v2

@@ 5.2 Main results
We start our empirical validation by comparing the proposed approach’s performance to existing state-of-the-art CLIP-based TTA approaches across the different datasets.
Evaluation in the presence of natural or no domain shift.  Table 1 reports the performance of the different methods for CIFAR-10, CIFAR-10.1, and CIFAR-100, which introduce a natural or no domain shift. The proposed method consistently outperforms existing approaches, even when the domain shift is small, highlighting the usefulness of resorting to our approach over CLIP zero-shot predictions.
@@
Impact of the performance under common corruptions.  Furthermore, in Table 1, we report a detailed analysis of the results obtained on CIFAR-10C and CIFAR-100C, which show the superiority of CLIP-OT in adapting CLIP in the presence of common corruptions. In particular, compared to the baseline CLIP, our model brings performance gains of nearly $18\%$ in CIFAR-10C and CIFAR-100C, without additional supervision. These performance gains are similar when compared to other popular TTA methods, e.g., TENT, or TPT, with differences ranging from $10\%$ to $21\%$. While this gap is reduced compared to very recent approaches, i.e., both versions of WATT and CLIPArTT, per-corruption classification accuracy differences are still significant, with CLIP-OT outperforming these two approaches by up to $5\%$ compared to WATT, i.e., $5.14\%$ in pixelate (CIFAR-10C) or $3.11\%$ in Glass Blur (CIFAR-100C), and nearly $11\%$ compared to CLIPArTT, e.g., $10.62\%$ in pixelate and $8.11\%$ in Gaussian Noise. As average, CLIP-OT improves the performance of very recent state-of-the-art approaches by a substantial difference of $3.24\%$ (CIFAR-10C), and $1.76\%$ in the more challenging scenario of CIFAR-100C, which contains $\times$10 number of classes.
What if the number of classes further increases?  In the following, we assess the performance of all methods on Tiny-ImageNet and Tiny-ImageNet-C benchmarks, which bring additional complexity due to the increased amount of categories present. From these results (Table 2), we can observe that the findings reported in previous sections hold, i.e., CLIP-OT significantly outperforming very recent state-of-the-art TTA approaches. More concretely, compared to WATT-S, the average improvement over all the corruptions is above $7.0\%$, with our method showing performance gains superior to $10\%$ in multiple corruptions, e.g., Shot Noise or Pixelate, and above $4.5\%$ in all remaining corruptions. Given the complexity and additional challenges that Tiny-ImageNet-C brings compared to CIFAR versions, as well as the non-negligible improvement provided by our method, we believe that these results showcase the clear superiority of CLIP-OT in more challenging scenarios.
Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-S NeurIPS’24 CLIP-OT Ours
Tiny-ImageNet 58.29 57.72 58.90 59.85 61.35 63.69$\blacktriangle$
Gaussian Noise 7.08 8.01 9.29 14.44 13.02 21.40 $\blacktriangle$
Shot Noise 9.41 10.04 11.70 17.44 15.94 24.90 $\blacktriangle$
Impulse Noise 3.44 4.18 4.85 10.37 6.90 17.34 $\blacktriangle$
Defocus Blur 21.71 24.53 27.56 31.46 29.91 35.39 $\blacktriangle$
Glass Blur 9.12 10.09 11.03 15.84 14.01 21.16 $\blacktriangle$
Motion Blur 34.52 36.94 38.97 41.34 41.26 46.26 $\blacktriangle$
Zoom Blur 27.44 29.48 34.29 35.06 33.96 40.93 $\blacktriangle$
Snow 32.51 32.20 34.45 36.86 37.76 42.32 $\blacktriangle$
Frost 36.33 35.72 37.13 38.20 39.65 44.97 $\blacktriangle$
Fog 25.94 27.46 28.89 33.44 32.13 38.60 $\blacktriangle$
Brightness 40.15 39.79 43.31 46.43 46.93 53.10 $\blacktriangle$
Contrast 1.81 2.24 3.15 6.24 3.53 11.88 $\blacktriangle$
Elastic Transf. 30.40 31.92 33.88 33.89 35.01 40.73 $\blacktriangle$
Pixelate 22.78 24.79 27.70 34.85 31.55 41.84 $\blacktriangle$
JPEG Compr. 29.59 30.93 33.60 37.32 36.46 42.86 $\blacktriangle$
Mean 22.14 23.22 25.32 28.88 27.87 34.91 $\blacktriangle$
Table 2: Performance comparison on Tiny-ImageNet and Tiny-ImageNet-C corruption benchmarks using ViT-B/32 backbone.
Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-P NeurIPS’24 WATT-S NeurIPS’24 CLIP-OT Ours $\Delta$
PACS 93.65 93.81 93.23 93.95 94.81 94.80 95.94 +1.14 $\blacktriangle$
OfficeHome 77.53 77.68 77.20 77.56 78.88 78.83 80.15 +1.32 $\blacktriangle$
VisDA-C (3D) 84.43 84.86 79.35 85.09 85.42 85.36 90.73 +5.37 $\blacktriangle$
VisDA-C (YT) 84.45 84.68 83.57 84.40 84.57 84.69 85.44 +0.75 $\blacktriangle$
Table 3: Performance comparison across PACS, OfficeHome, and VisDA-C datasets using ViT-B/32 backbone. $\Delta$ represents the difference between CLIP-OT and the State-of-the-art WATT-S [31]. Extended numerical values are presented in Appendix C.2, Table 9.
Performance under simulated and video shifts.  Table 3 (VisDA-C) reports the results of simulated (3D) and video (YT) shifts, which showcase the superiority of CLIP-OT compared to other methods. More concretely, the average difference compared to SoTA (e.g., WATT-S) is $3.03\%$, whereas per-dataset differences can go up to $11\%$, e.g., $11.38\%$ compared to TPT in 3D domain.
Texture and style shifts.  We now evaluate CLIP-OT on different distributional drifts, which may bring additional complexity. The results from these experiments (Table 3) show that the trend of superior performance is observed in almost all datasets (i.e., PACS and OfficeHome).
5.3 Ablation studies
@@
Figure 2: Ablation on each component across multiple datasets. Different results come from adding each element of the proposed formulation. In green, performance differences compared to CLIP. The numerical values are provided in Appendix C.
Impact of each component.  In Figure 2, we empirically validate the benefit of each component in CLIP-OT design. Each configuration is detailed in Appendix C.1. First, we can observe that computing the code $Q^{*}_{m}$ from each template $m$ (Eq. (9)) and averaging those to obtain the final predictions consistently improves over zero-shot CLIP. It is noteworthy to highlight that the transportation polytope $\mathcal{Q}$ enforces additional explicit constraints on $\mathbf{Q}$, which ensure a well-defined probabilistic assignment and “break” the direct dependency between $\mathbf{Q}$ and $\mathbf{P}$. This explains the differences between direct CLIP predictions and this version of our approach, which directly leverages $\mathbf{Q}$ as predictions. Furthermore, while this performance is far from state-of-the-art approaches, this strategy is training-free, not requiring additional computational complexity due to gradient descent updates. Next, Average Template employs the pseudo-codes computed in the previous version to minimize the cross-entropy in Eq. (3). Furthermore, the prediction for the $i$-th image, $\mathbf{p}_{i}$, is computed with the average template for the class $k$, i.e., $\mathbf{t}_{k}$. This shows that leveraging the obtained pseudo-codes to refine the model leads to a performance improvement, which is consistent across all datasets. Last, we introduce the novel knowledge distillation strategy proposed in this work, where distinct templates are used to get multiple pseudo-codes and guide the model training, i.e., CLIP-OT. Its substantial performance gains are especially notable in the most challenging scenarios, e.g., CIFAR-100C.
On the impact of multiple templates.  We now assess how the performance evolves as the number of templates increases (Table 4). Concretely, we trained our model multiple times, varying the number of templates $N\in\{1,\dots,8\}$. For each $N$, the average accuracy over all template combinations was computed, leading to 256 experiments per dataset. Results from this experiment show that CLIP-OT performance consistently increases with the number of templates. This showcases how leveraging multiple templates individually can extract richer information to adapt CLIP at test time.
##\## Templates 1 2 3 4 5 6 7 8
CIFAR-10 91.59 92.16 92.46 92.69 92.88 93.04 93.14 93.14
CIFAR-100 68.36 69.46 70.08 70.56 70.90 71.15 71.38 71.34
Tiny-ImageNet 60.62 61.73 62.32 62.69 62.95 63.16 63.20 63.11
Table 4: Accuracy (%) of CLIP-OT with varying numbers of templates (see Appendix A) on natural domain shift benchmarks.
@@
Do observations hold across backbones?  Table 5 reports the performance across different datasets when ViT-B/16 is employed as the visual encoder of CLIP. These results demonstrate that CLIP-OT is model agnostic, as its superiority is consistently maintained across different backbones.
Dataset CLIP TENT TPT CLIPArTT WATT-S CLIP-OT
CIFAR-10 89.25 92.75 89.90 92.61 91.97 94.69
CIFAR-10.1 84.00 88.52 83.75 88.72 88.10 89.25
CIFAR-10C 60.15 68.00 59.75 73.22 76.22 80.11
CIFAR-100 64.76 71.73 67.15 71.34 72.85 74.06
CIFAR-100C 32.01 37.90 33.73 40.08 48.95 51.24
TinyImageNet 58.12 63.59 63.87 64.73 66.05 66.34
TinyImageNet-C 20.92 29.78 26.96 32.90 31.66 37.69
Table 5: Quantitative results with ViT-B/16 as backbone for TTA across different datasets and baselines.
@@
Computational analysis.  Figure 3 depicts the running time required for relevant baselines and the proposed CLIP-OT across several datasets, ordered in terms of increasing complexity, i.e., number of categories. These results expose that different TTA methods, particularly SoTA, substantially differ in total runtimes. In particular, recent CLIPArTT and WATT-S constantly increase the required runtime with the number of classes, driven by their iterative nature on the multiple template evaluations and weight averaging during inference. In contrast, CLIP-OT avoids this overhead by distilling this information during the adaptation stage and computing multiple text embeddings off-line only once. Moreover, the Sinkhorn algorithm—used for generating pseudo-codes is highly efficient, as its iterative updates converge faster on the low-dimensional similarity matrices and benefit from parallelizable computations, accounting for only nearly $1\%$ of the total runtime. This combination of speed and consistency establishes CLIP-OT as a scalable and effective method, providing runtimes from one to two orders of magnitude lower than CLIPArTT and WATT-S for the most complex dataset (i.e., Tiny-ImageNet), while significantly outperforming them from a discriminative standpoint.
Figure 3: Computational efficiency. The runtime of recent CLIP TTA methods is assessed, using a standard input batch of size 128. Runtime differences compared to our CLIP-OT highlighted in red.

### v3

@@ 5.2 Main results
Performance under common visual corruptions. We present a unified evaluation across four standard benchmarks in Table 1, which shows the superiority of SAT in adapting CLIP in the presence of common corruptions against a comprehensive suite of recent TTA methods. Compared to vanilla CLIP, our model brings performance gains of $17.8\%$ (CIFAR-10C) and $17.9\%$ (CIFAR-100C) without requiring additional supervision. These performance gains are similar when compared to other popular TTA methods, e.g., TENT or TPT, with differences ranging from $10\%$ to $21\%$. While this gap is reduced compared to recent approaches, such as WATT, CLIPArTT, and BATCLIP, the differences remain significant. SAT outperforms the second-best competitor, i.e., WATT, by up to $3.2\%$ in CIFAR-10C (e.g., $5.1\%$ in pixelate) and $1.8\%$ in the more challenging scenario of CIFAR-100C (which contains $\times$10 classes). Additionally, it achieves significant gains over recent baselines in specific corruptions, e.g., $19.8\%$ compared to BATCLIP in pixelate (CIFAR-10C) or $16.3\%$ in Glass Blur (CIFAR-100C). Also, compared to CLIPArTT, SAT shows gains of nearly $10.6\%$ on pixelate (CIFAR-100C) and $8.1\%$ on Gaussian Noise (CIFAR-100C).
Table 1: Performance in visual corruptions benchmarks. Results using CLIP ViT-B/32. Best method in bold, second best underlined.
Method
Gaussian
Shot
Impulse
Defocus
Glass
Motion
Zoom
Snow
Frost
Fog
Bright
Contrast
Elastic
Pixelate
JPEG
Mean $\uparrow$
CIFAR-10C
CLIP [ICLR’21] 35.27 39.67 42.61 69.76 42.40 63.97 69.83 71.78 72.86 67.04 81.87 64.37 60.83 50.53 55.48 59.22
TENT [ICLR’21] 41.27 47.20 48.58 77.12 52.65 71.25 76.20 78.29 79.84 77.39 87.78 79.47 70.00 63.74 62.64 67.56
SAR [ICLR’22] 47.58 50.39 47.19 71.65 49.34 70.27 72.63 71.66 72.82 69.48 82.34 70.54 60.98 48.07 58.48 62.89
VTE [ECCVw’24] 44.40 47.70 42.90 64.90 45.00 66.70 67.00 67.40 64.50 65.30 74.90 53.60 61.20 42.60 50.80 57.26
TPT [NeurIPS’22] 33.90 38.20 37.66 67.83 38.81 63.39 68.95 70.16 72.39 64.31 81.30 62.26 56.43 42.80 53.67 56.80
WATT [NeurIPS’24] 63.84 65.28 58.64 78.94 65.12 77.81 79.32 79.79 80.54 78.53 87.11 81.20 72.66 71.11 67.36 73.82
CLIPArTT [WACV’25] 59.90 62.77 56.02 76.74 61.77 76.01 77.40 77.29 79.20 75.74 86.59 77.82 70.20 66.52 63.51 71.17
BATCLIP [ICCV’25] 50.89 56.01 54.35 76.17 56.11 74.71 76.01 77.74 79.33 75.87 86.46 78.65 68.76 56.41 61.79 68.62
SAT (Ours) 64.85 67.34 62.27 82.09 68.07 81.30 83.13 83.71 83.40 82.56 89.90 84.86 76.08 76.25 70.03 77.06
CIFAR-100C
CLIP [ICLR’21] 14.80 16.03 13.85 36.74 14.19 36.14 40.24 38.95 40.56 38.00 48.18 29.53 26.33 21.98 25.91 29.43
TENT [ICLR’21] 14.38 17.34 10.03 49.05 3.71 46.62 51.84 46.71 44.90 47.31 60.58 45.90 33.09 26.47 29.89 35.19
SAR [ICLR’22] 22.82 25.10 18.68 44.51 21.78 43.04 47.04 46.75 47.34 44.62 57.00 42.17 31.51 25.09 30.83 36.55
VTE [ECCVw’24] 10.00 10.30 13.30 36.10 20.40 37.90 39.80 42.20 40.80 36.60 45.50 29.20 30.80 17.00 20.70 28.71
TPT [NeurIPS’22] 14.03 15.25 13.01 37.60 16.41 37.52 42.99 42.35 43.31 38.81 50.23 28.09 28.12 20.43 28.82 30.46
WATT [NeurIPS’24] 32.07 34.36 30.33 52.99 32.15 50.53 55.30 52.77 53.79 51.49 63.57 52.76 40.90 40.97 39.59 45.57
CLIPArTT [WACV’25] 25.32 27.90 25.62 49.88 27.89 47.93 52.70 49.72 49.63 48.77 61.27 48.55 37.45 33.88 36.07 41.51
BATCLIP [ICCV’25] 17.25 19.76 18.98 42.20 19.00 40.81 46.59 41.34 40.14 41.56 53.85 34.07 31.38 25.51 28.96 33.43
SAT (Ours) 33.43 35.60 30.94 53.87 35.26 52.77 56.71 54.30 54.92 53.57 64.43 55.01 43.79 44.51 40.83 47.33
Tiny-ImageNet-C
CLIP [ICLR’21] 7.08 9.41 3.44 21.71 9.12 34.52 27.44 32.51 36.33 25.94 40.15 1.81 30.40 22.78 29.59 22.15
TENT [ICLR’21] 8.01 10.04 4.18 24.53 10.09 36.94 29.48 32.20 35.72 27.46 39.79 2.24 31.92 24.79 30.93 23.22
SAR [ICLR’22] 9.09 10.94 3.65 5.50 1.68 14.02 12.08 20.72 24.62 8.37 32.35 0.71 15.32 12.39 25.35 13.12
VTE [ECCVw’24] 18.63 20.34 4.71 9.62 2.21 30.37 21.68 38.84 40.27 17.41 41.22 0.63 31.64 25.33 37.79 22.71
TPT [NeurIPS’22] 9.29 11.70 4.85 27.56 11.03 38.97 34.29 34.45 37.13 28.89 43.31 3.15 33.88 27.70 33.60 25.32
WATT [NeurIPS’24] 13.02 15.94 6.90 29.91 14.01 41.26 33.96 37.76 39.65 32.13 46.93 3.53 35.01 31.55 36.46 27.87
CLIPArTT [WACV’25] 14.44 17.44 10.37 31.46 15.84 41.34 35.06 36.86 38.20 33.44 46.43 6.24 33.89 34.85 37.32 28.88
BATCLIP [ICCV’25] 11.96 15.48 10.05 31.89 14.76 43.31 39.07 39.02 39.05 31.91 49.06 5.65 32.79 36.63 39.12 29.32
SAT (Ours) 21.40 24.90 17.34 35.39 21.16 46.26 40.93 42.32 44.97 38.60 53.10 11.88 40.73 41.84 42.86 34.91
ImageNet-C
CLIP [ICLR’21] 11.30 11.58 12.28 20.88 8.92 19.78 17.62 19.92 23.48 25.90 47.34 15.48 17.02 28.00 27.60 20.47
TENT [ICLR’21] 8.00 7.20 9.20 23.04 10.84 22.86 19.04 21.24 23.86 26.54 48.54 18.32 17.90 30.32 29.66 21.10
SAR [ICLR’22] 13.07 15.69 13.92 22.74 14.53 23.41 19.49 22.65 24.89 29.47 48.39 18.88 19.61 31.68 29.07 23.17
VTE [ECCVw’24] 7.12 10.24 9.18 27.31 10.27 26.42 27.36 24.28 26.15 31.22 49.37 13.09 14.18 32.44 31.33 22.66
TPT [NeurIPS’22] 8.94 7.22 7.55 20.47 9.13 21.78 23.92 24.61 21.54 24.98 40.37 15.22 13.18 30.74 24.63 20.01
WATT [NeurIPS’24] 7.76 7.06 8.94 24.16 12.46 25.00 21.52 21.58 24.16 26.62 49.74 21.14 19.90 32.70 32.16 22.33
CLIPArTT [WACV’25] 14.74 15.10 15.30 10.82 9.02 13.82 12.30 16.96 22.52 19.90 41.78 0.26 12.84 22.80 31.94 18.15
BATCLIP [ICCV’25] 14.84 15.10 15.52 24.42 17.18 25.64 23.08 25.06 25.58 31.08 49.66 18.44 22.20 33.42 33.02 24.95
SAT (Ours) 18.98 24.54 25.04 25.46 21.98 26.34 24.46 27.78 28.74 29.32 41.98 17.90 25.70 26.32 25.68 26.01
What if the number of classes increases? The gap is even more pronounced on the more difficult Tiny-ImageNet-C testbed, where our $34.9\%$ mean accuracy is $5.6\%$ higher than that of the strongest baseline, BATCLIP. Finally, on the highly challenging 1000-class ImageNet-C benchmark, where many methods show limited gains, SAT again proves its robustness. Specifically, achieves the highest mean accuracy, $26.0\%$, a promising $5.5\%$ gain over the CLIP baseline, outperforming all other TTA baselines.
Table 2: Performance comparison under domain shifts. Results using CLIP ViT-B/32. Best method in bold, second best underlined.
Method PACS O-Home VLCS VisDA-3D VisDA-YT Mean
(texture / style) (simulated / video)
CLIP 93.65 77.53 80.16 84.43 84.45 84.04
TENT 93.81 77.68 80.27 84.86 84.68 84.26
TPT 93.23 77.20 74.57 79.35 83.57 81.58
CLIPArTT 93.95 77.56 80.06 85.09 84.40 84.21
WATT 94.80 78.83 81.14 85.36 84.69 84.96
BATCLIP 94.52 78.90 80.78 81.97 83.60 83.95
SAT (Ours) 95.94 80.15 78.33 90.73 85.44 86.12
Performance under additional shifts:  Table 2 (VisDA-C) reports the results of simulated (3D) and video (YT) shifts, which showcase the superiority of SAT. More concretely, the differences compared to the second-competitor SoTA method (WATT) in these two datasets are $5.4\%$ and $0.8\%$. Regarding other popular baselines, per-dataset differences can go up to $11\%$, e.g., $11.4\%$ compared to TPT in the 3D domain. It is worth noting also the significant performance drop of the very recent BATCLIP on this benchmark, which underscores the instability of methods that rely too heavily on their own initial predictions in the face of large domain gaps. Furthermore, Table 2 results on texture and style shifts show that the trend of superior performance is observed in other datasets (i.e., PACS and OfficeHome), whereas performance is degraded in only one out of the five datasets, i.e., VLCS. All considered, SAT emerges as a powerful and reliable TTA approach, where average improvement gains compared to recent methods are $2.2\%$ (BATCLIP), $1.2\%$ (WATT), and $1.9\%$ (CLIPArTT), showcasing its versatility.
5.3 In-depth studies
Impact of each component. In Fig. 3, we empirically validate the benefits of each component in SAT (each configuration is detailed in Appendix B.1). The ‘Training-Free OT’ (using OT for assignments but not updating the model) already improves over zero-shot CLIP, demonstrating the power of our batch-aware assignment. ‘Average Template’ (using an averaged text prototype) provides a further boost. Finally, our full method, which utilizes ‘Multi-Templates’ knowledge distillation, yields the largest gains, demonstrating that all components are necessary. The benefit of leveraging multiple textual semantic views is also supported in Appendix B.5 experiments, where we showcase that performance consistently scales with the number of templates.
@@ 5.3 In-depth studies
Figure 3: Ablation on each component. Results come from adding each element of SAT. In green, performance differences compared to CLIP (numerical values in Appendix B.2).
@@
Do observations hold across backbones? Tab. 3 reports the performance across different datasets when larger CLIP pre-trained models are employed. These results demonstrate the model-agnostic nature of SAT, as its superiority remains consistent across different backbones.
Table 3: TTA results using larger CLIP ViT backbones.
a) ViT-B/16 CIFAR-10C CIFAR-100C Tiny-IN-C ImageNet-C
CLIP 60.15 32.01 20.92 20.89
TENT 68.00 37.90 29.78 22.79
TPT 59.75 33.73 26.96 22.35
CLIPArTT 73.22 40.08 32.90 23.47
WATT 76.22 48.95 31.66 24.38
BATCLIP 73.52 38.85 29.30 27.16
SAT 80.11 ($\uparrow$ 19.96) 51.24 ($\uparrow$ 19.23) 37.69 ($\uparrow$ 16.77) 28.35 ($\uparrow$ 7.46)
b) ViT-L/14 CIFAR-10C CIFAR-100C Tiny-IN-C ImageNet-C
CLIP 76.04 44.59 34.98 32.05
TENT 79.18 50.14 40.28 33.09
TPT 75.01 47.58 41.07 30.67
CLIPArTT 78.06 52.52 42.98 34.13
WATT 80.06 54.34 43.28 36.30
BATCLIP 83.79 48.84 35.72 37.07
SAT (Ours) 86.35 ($\uparrow$ 10.31) 62.21 ($\uparrow$ 17.62) 50.36 ($\uparrow$ 15.38) 38.82 ($\uparrow$ 6.77)
Generalization across VLMs. SAT’s principles are also model-agnostic beyond the CLIP pre-training framework. As shown in Table 4, SAT consistently delivers state-of-the-art performance when applied to a modern SigLIP backbone. The gains are even more pronounced on this architecture, with a $15.0\%$ improvement on CIFAR-100C and $5.4\%$ on ImageNet-C over zero-shot SigLIP, proving the general applicability of our cross-modal alignment paradigm.
Table 4: Performance using SigLIP [zhai2023sigmoidlosslanguageimage].
Method CIFAR-10C CIFAR-100C Tiny-IN-C ImageNet-C
SigLIP 59.04 34.76 22.00 26.49
BATCLIP 67.45 39.26 24.72 30.40
SAT (Ours) 76.20 ($\uparrow$ 17.16) 49.72 ($\uparrow$ 14.96) 26.81 ($\uparrow$ 4.81) 31.92 ($\uparrow$ 5.43)
@@
Computational analysis. Fig. 4 depicts the running time required for relevant baselines and the proposed SAT across several datasets, ordered in terms of increasing complexity, i.e., number of categories. These results expose that different TTA methods, particularly SoTA, substantially differ in total runtimes. In particular, recent CLIPArTT and WATT constantly increase the required runtime with the number of classes, driven by their iterative nature, which involves multiple template evaluations and weight averaging during inference. BATCLIP is also highly efficient, yet its runtime still shows a slight increase as the number of classes grows. In contrast, SAT avoids this overhead by distilling this information during the adaptation stage and computing multiple text embeddings off-line only once. It is worth noting that the Sinkhorn algorithm used for generating pseudo-codes is highly efficient, as its iterative updates converge faster on the low-dimensional similarity matrices and benefit from parallelizable computations, accounting for only nearly $1\%$ of the total runtime. This combination of speed and consistency establishes SAT as a scalable and effective method, providing runtimes from one to two orders of magnitude lower than CLIPArTT and WATT for the most complex dataset (i.e., Tiny-ImageNet), while significantly outperforming them from a discriminative standpoint.
Figure 4: Inference Runtime Comparison. Runtime in seconds for ViT-B/32 TTA methods on an NVIDIA RTX A6000 using a test-time batch of size=128 images.